OpenStudy (anonymous):

Lucy is planning to run for an hour, and she's using her iPod for the countdown, so she has to create a 60-minute playlist. Her iPod contains: six 5-min songs; five 6-min songs; four 8-min songs. Count the number of possibles playlists where the order doesn't matter.

OpenStudy (kittiwitti1):

\[60 \div 8 = 7.5\text{ 8 minute songs}\]

OpenStudy (kittiwitti1):

\[\text{This is only one possible playlist, btw}\] We only have four 8 minute songs, so now we need a value that makes up the rest (3.5 eight-minute songs)

OpenStudy (anonymous):

I guess that's not kind of right

OpenStudy (kittiwitti1):

I'm working on actually getting all the results. \[\text{A mafia boss must be patient. ;) }\]

OpenStudy (anonymous):

Okayyy :)

OpenStudy (kittiwitti1):

(I forgot how to do this so I asked someone else I'm sorry ; o; )

OpenStudy (jhannybean):

Seems like a combination/permutation like problem.

OpenStudy (kittiwitti1):

I was thinking that too, but I suck at those. ; o; @Jhannybean

ganeshie8 (ganeshie8):

Say she puts \(a\) number of 5-min songs, \(b\) number of 6-min songs and \(c\) number of 8-min songs in her playlist. then we need to solve below equaiton in integers : \[5a+6b+8c = 60\] \(0\le a\le 6\) \(0\le b\le 5\) \(0\le c\le 4\)

ganeshie8 (ganeshie8):

Notice \(a\) cannot be an odd number because that evaluates the left hand side to an odd number. so \(a\) can only take values : \(\{0,2,4,6\}\)

OpenStudy (kittiwitti1):

Thank you for saving the day @ganeshie8 ! T ^T

OpenStudy (anonymous):

Wow never thought of that. Then what could have been the next? Sorry late reply